OpenStudy (anonymous):

If the first term of an arithmetic sequence is 4 and the third term is 18, what is the 23rd term?

OpenStudy (anonymous):

Help Please!

OpenStudy (anonymous):

Hi! What have you tried so far? (Or are you completely stuck?)

OpenStudy (anonymous):

ive trying to do 4+923-1)(18), but its wrong

OpenStudy (anonymous):

4+(23-1)(18)

OpenStudy (anonymous):

Okay, that's a good start. First of all, if the first term is 4 and the third term is 18, what do you think the second term is?

OpenStudy (anonymous):

im not really sure how to figure out the second term that well

OpenStudy (anonymous):

Okay, that's no problem! In general, the first term of a series will be A, the second one will be A + D, the third will be A + 2D and so on, so the (n)th term will be A + (n-1)D. You might have used different letters instead of A and D, but I'm guessing you've seen that formula before?

OpenStudy (anonymous):

yeah the formula that i was given was an=a1+(n-1)(d)

OpenStudy (anonymous):

Great! In that case, in your question, you're told that the first term is 4, so A = 4, and the third term is 18, so A + 2D = 18. Using that, can you find what D is?

OpenStudy (anonymous):

Or would i have to do, an=2(n)-4

OpenStudy (anonymous):

Hmm, not sure about that. I'd suggest trying to find D - can you see how?

OpenStudy (anonymous):

how do i figure out the second term from the 1st and 3rd

OpenStudy (anonymous):

Once you know A and D, you can figure out any term that you like.

OpenStudy (anonymous):

(Using your formula).

OpenStudy (anonymous):

Just a formula, any formula

OpenStudy (anonymous):

The formula you were given was: \[a_{n} = a_{1} + (n-1) d\]

OpenStudy (anonymous):

18=4+(3-1)d

OpenStudy (anonymous):

yeah basketweave

OpenStudy (anonymous):

find out d...u can easily find out the 23rd term.

OpenStudy (anonymous):

Okay, so you know that a1 is 4, and a3 is 18, right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So using your formula, with n=3, we get\[a_{3}=a_{1}+(3-1)d\]

OpenStudy (anonymous):

Can you see how you might find d using that?

OpenStudy (anonymous):

(It's ok if you're unsure!)

OpenStudy (anonymous):

It's got nothing to do with a2.

OpenStudy (anonymous):

that cannot be a2

OpenStudy (anonymous):

i need the second term then i will get the difference but i have no clue how to find it

OpenStudy (anonymous):

Okay, so a3-a2 = d, and a2-a1=d, right?

OpenStudy (anonymous):

If the first term of an arithmetic sequence is 4 and the third term is 18, what is the 23rd term? A. 92 B. 116 C. 158 D. 161 These are the answers

OpenStudy (anonymous):

it should be a3..in ur formula becase u taken n=3... @BasketWeave n u got ur d..add d with first term you got ur 2nd term...

OpenStudy (anonymous):

d=7,a=4 find out 23rd term

OpenStudy (anonymous):

thanks i can get it form there

OpenStudy (anonymous):

I got c

OpenStudy (anonymous):

Maybe a picture will help: